OpenStudy (anonymous):

The solutions to a certain quadratic equation are x = -4 and x = 3. Write the equation in standard form below.

OpenStudy (er.mohd.amir):

if roots are x1,x2 then its sum=x1+x2 and product=x1*x2 then equation is x^2-(sum)*x+product=0 hence eq is x^2+x-12=0

OpenStudy (anonymous):

thanks!